OpenStudy (anonymous):

Note that in an isosceles triangle, two angles have the same measure and the sum of all three angle measures is 180 degrees. Set up and equation and solve: Determine the measure (in degrees) of the third angle in an isosceles triangle, if the third angle is 10 more than one-half of either of the equal measure angles.

OpenStudy (anonymous):

since two angles are equal let's say they both equal x. ok? so angle 1 =x and angle 2 =x since angle 1 and 2 are the same. not it says the third angle is 10 more than one half of either one. so third angle = 10 + (x/2) so now you can add all the three angles up and make them equal to 180 and you will find your answer. so x + x + 10 + (x/2) = 2x + 10 + (x/2) = 5x/2 + 10 = 180 so 5x/2 = 170 5x = 340 x = 340/5 x = 68

OpenStudy (anonymous):

so angle 1 = 68 angle 2 = 68 angle 3 = 10+(68/2) = 10 + 34 = 44 and if you add all those up they will equal to 180